OpenStudy (anonymous):

Simplify 5^3. 15 125 225

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

well 5 raised to the pwer of 5 is 25

OpenStudy (anonymous):

i think the answer is 125

OpenStudy (anonymous):

i mean 5 raised to the power of 3

OpenStudy (ghostedly):

5^3 =5*5*5 =53

OpenStudy (anonymous):

dude no 5*5=25 25*5=125

OpenStudy (anonymous):

umm.........................................

OpenStudy (anonymous):

put it in a calc

OpenStudy (ghostedly):

I'm using my head xD

OpenStudy (zarkon):

these posts are making me sad

OpenStudy (anonymous):

lol

OpenStudy (ghostedly):

But answer is 125

OpenStudy (anonymous):

its exponents

OpenStudy (anonymous):

hallo zarkon

OpenStudy (anonymous):

as my mother used to always say "If you dont have anything good to say shut your mouth"

TheSmartOne (thesmartone):

Same Zarkon. Such easy questions, and all these direct answers.